# Redefining the Shortest Path Problem Formulation of the Linear Non-Gaussian Acyclic Model: Pairwise Likelihood Ratios, Prior Knowledge, and Path Enumeration

### v1

@@ 4.1.2. Comparing PLR LiNGAM-SPP with Existing Methods
Furthermore, Table 2 presents the average run times for the same cases. As expected, DirectLiNGAM, being a greedy algorithm, exhibits the fastest runtimes. Meanwhile, the runtimes of copent and $k$NN LiNGAM-SPP, both of which rely on the $k$NN mutual information estimator, tend to increase with sample size ($N$). Despite this trend, copent LiNGAM-SPP has the smallest runtimes out of all LiNGAM-SPP methods. On the other hand, $k$NN LiNGAM-SPP runs notably slower; for instance, at $p=15$, the runtime can extend to hours. However, while copent LiNGAM-SPP is faster, its performance remains subpar. Overall, PLR LiNGAM-SPP presents the most favorable compromise between runtime efficiency and performance, with the added advantage of not requiring parameter tuning.
4.1.3. Testing on Real-world Data
To evaluate PLR LiNGAM-SPP on real-world data, we sourced datasets from the CMU-CLeaR (Carnegie Melon University Causal Learning and Reasoning Group) list of benchmarks, accessible on GitHub: https://github.com/cmu-phil/example-causal-datasets. Our criteria for selection included datasets with continuous features and known ground truth. Additionally, we restricted our choices to datasets with $p\leq 20$ due to the computational demands of the LiNGAM-SPP method for larger values of $p$.
Table 3. Performance Comparison of PLR DirectLiNGAM and PLR LiNGAM-SPP on Real-World Datasets
Known Edges DirectLiNGAM LiNGAM-MMI
@@ 4.1.3. Testing on Real-world Data
To derive the causal graph from LiNGAM-SPP, we adapted code from the lingam package (Ikeuchi et al., 2023). Specifically, after getting the causal order, we used the same sparse regression as DirectLiNGAM to identify which edges are zero (Zou, 2006; Shimizu et al., 2011). Table 3 provides the ground truth for each dataset, indicating required and forbidden edges, along with the edges captured by PLR DirectLiNGAM and PLR LiNGAM-SPP. The ground truth originally included tiers, which we converted into forbidden edges, given that later tiers cannot cause earlier ones. For more details about the datasets and the provided ground truth, please refer to the GitHub repository.
The results in Table 3 reveal that for the wine-quality (red and white) datasets, PLR LiNGAM-SPP outperforms PLR DirectLiNGAM, capturing fewer forbidden edges and more required edges. For the other datasets, both methods yield identical results. This observation aligns with the findings in Table 1, indicating that PLR LiNGAM-SPP is more robust for longer sequences. Moreover, this suggests that, in most cases, PLR LiNGAM-SPP performs at least as well as PLR DirectLiNGAM.
4.2. Incorporating Known Relative Ordering
@@
We demonstrate the impact of incorporating prior knowledge on performance in Figure 2 and Table 4. To quantify the extent of prior knowledge, we express it as a percentage of variables with specified orderings. For example, if $p=8$, 50% knowledge means that the relative ordering of 4 variables is given. In this demonstration, the input is the relative ordering of 2 or more variables, which we then convert to variable pairs, e.g., $(1,2,3)$ implies $(1,2)$, $(2,3)$, $(1,3)$. However, it is also possible to manually specify pairs. As expected, incorporating more prior knowledge improves performance (Figure 2) and generally reduces the traversed number of edges (Table 4).
Relative ordering provides a more flexible way of incorporating prior knowledge, which is valuable because we do not always know the exact edges. It is more common for us to know the relative ordering rather than the exact edges. For instance, in most applications, it is known that nothing can affect a person’s age, sex, and other background variables, so these must be positioned earlier in the ordering relative to other variables. Another example is incorporating temporal information, where features with time-related attributes may exist, and we only know the temporal order but not necessarily the ordering among the features within the same time slice.
4.3. Predicting Causal Graph Properties from Path Distributions
@@
Our motivation for predicting causal graph properties from path distributions is drawn from the observations depicted in Figure 3. These histograms show the combined log-transformed PLR path distributions for $p=4,5,6$ (where $p$ represents the number of features), using the same data generating process and simulation configurations as detailed in preceding sections. Notably, discernible patterns emerge in the path distributions across different scenarios: More complex scenarios, characterized by the presence of confounders or higher graph sparsity, exhibit an increased frequency of shorter paths. This observation suggests that determining the correct order becomes more challenging and ambiguous under such conditions.
4.3.1. Confounder Detector
Detecting unmeasured confounders is crucial, especially for LiNGAM, which assume the absence of unmeasured confounders. In this section, we developed a confounder detector—a classifier for detecting the presence of unmeasured confounders. The features used were the 3rd to the 30th standard moments of the log-transformed path distribution, and the target is binary: 1 if there are confounders and 0 if none. We then trained different classifier models, as shown in Table 5. Note that, for demonstration purposes, these models were not fine-tuned, so there is still room for improving these results. Regardless, the models demonstrate good performance, with CatBoost achieving the highest AUC of $0.78$, indicating good discrimination between the classes.
Table 5. Comparison of Confounder Detectors Trained with Various Classifiers
@@ 4.3.1. Confounder Detector
Figure 4 shows the ROC curve and the confusion matrix from CatBoost evaluated on the test set. The optimal threshold is determined using Youden’s J statistic, which selects the threshold that maximizes the difference between the true positive and false positive rates. This threshold can be changed depending on the precision and sensitivity requirements of the use case. However, for the optimal threshold, we see that the CatBoost model exhibited good sensitivity (77.4% recall) and precision (69.6%).
@@
The figure illustrates the Receiver Operating Characteristic (ROC) curve and confusion matrix for the confounder detector model. Employing CatBoost and assessed on the test set, the optimal threshold, identified through Youden’s J statistic, maximizes the difference between true positive and false positive rates. The associated confusion matrix is derived from this threshold, showcasing the CatBoost model’s remarkable sensitivity (77.4% recall) and precision (69.6%).
Thus, we have demonstrated that it is possible to detect unmeasured confounders with reasonable performance using information from the shape of the path distribution captured through the standardized moments.
4.3.2. Sparsity Estimator
Another task we explore is predicting graph sparsity using standard moments. Using the same features as the confounder detector, we initially create a classifier predicting whether sparsity is greater than 0.5. Table 6 presents the results from different models. Similar to the confounder detector, CatBoost achieved the highest AUC of 0.77. While CatBoost’s precision and recall are remarkable, AdaBoost offers slightly better precision, and XGBoost provides slightly better recall—choosing between them depends on the specific use case. Figure 5 displays the ROC curve and confusion matrix for CatBoost. Once again, it demonstrates good AUC, precision, and recall, indicating the model’s effective discrimination.
Table 6. Comparison of Sparsity Estimators Trained with Various Classifiers
@@ 4.3.2. Sparsity Estimator
The figure illustrates the Receiver Operating Characteristic (ROC) curve and confusion matrix for the sparsity estimator model. Employing CatBoost and assessed on the test set, the optimal threshold, identified through Youden’s J statistic, maximizes the difference between true positive and false positive rates. The associated confusion matrix is derived from this threshold, showcasing the CatBoost model’s remarkable sensitivity (72.17% recall) and precision (70.43%).
In addition to the classifier, we explored creating a regression model using the same features, with the target being the actual sparsity (ranging from 0 to 1). Table 7 compares results from different models. Among them, AdaBoost delivered the best coefficient of determination $R^{2}$ of $0.16$, indicating that the shape of the distribution (i.e., the standard moments) can explain some of the variance in sparsity. AdaBoost also yields an RMSE of $0.27$, reasonably estimating sparsity.
Table 7. Regression Results for Graph Sparsity Prediction
@@
4.3.3. Performance Predictors
Thus far, we have created models for detecting confounders and graph sparsity. These two properties are crucial for evaluating the performance of DirectLiNGAM and LiNGAM-SPP, as cases involving confounders or sparse graphs tend to pose greater challenges and lead to poorer performance for these methods. However, we wanted to test if it were possible to directly predict the performance of these methods, so we developed classifiers for PLR DirectLiNGAM and PLR LiNGAM-SPP to predict whether they would produce the correct causal ordering. In other words, the target is 1 if the predicted causal order is entirely correct. Table 8 showcases the results.
Table 8. Comparison of Performance Predictors Trained with Various Classifiers
@@ 4.3.3. Performance Predictors
For the LiNGAM-SPP performance predictor, CatBoost delivered the best AUC of 0.82. However, given the relatively small number of instances with exact correct ordering (as it is more likely to get some of the ordering wrong than to get everything right), while sensitivity (recall) is reasonable at  0.70, precision is not very high ($0.33$), indicating a significant number of false positives.
Meanwhile, for the DirectLiNGAM performance predictor, both Random Forest and CatBoost provided similar AUCs of $0.82$, with Random Forest having the best recall at 0.80, while another model, XGBoost, exhibited the highest precision at 0.31. Similar to the LiNGAM-SPP case, sensitivity of the models is good, but precision is not as high, suggesting a notable presence of false positives.
@@
We also explored fitting regression models to the ordering score of LiNGAM-SPP and DirectLiNGAM, with the results shown in Table 9. For the LiNGAM-SPP case, AdaBoost emerged as the best model, achieving the highest $R^{2}$ of  0.21, indicating its ability to explain some variance in the ordering score. AdaBoost also boasted the lowest RMSE of  0.18, reasonably pinpointing the ordering score within a reasonable range. Similarly, for DirectLiNGAM, AdaBoost also outperformed other models, achieving an $R^{2}$ of $0.21$ and an RMSE of $0.18$.
4.3.4. Path Sampling
One drawback of path enumeration is the factorial increase of the number of edges with the number of features. To address the impracticality of using the entire path distribution (especially in cases with large $p$), we opt for a more practical solution—sampling from the path distribution to compute the standardized moments. As mentioned earlier, ZDDs provide this functionality by sampling from the subset of graphs it represents. Here, we demonstrate the tradeoff between path sample size and model performance (AUC), revealing that sufficiently good performance for the models discussed earlier can still be achieved without resorting to the entire path distribution.
@@ 4.3.4. Path Sampling
Table 10 presents the performance of the models under various path sampling sizes. In this demonstration, we limit the training and testing to $p=7$ since $7!=5040$, which is sufficiently large for our evaluation. The table illustrates that the models’ AUC generally increases with the path sample size, approaching the performance of the full path distribution. For instance, even at a path sample size of 2500, the confounder detector and sparsity estimator already achieved AUCs that are sufficiently close to those obtained using the full path distribution. While the performance predictors for the same path sample size still have a slightly larger gap, the performance is already acceptable. This demonstrates the trade-off between path sample size and model performance. For larger values of $p$, even a fraction of the possible paths may already yield satisfactory performance.
@@ 5. Conclusion
In this study, we addressed a critical limitation of LiNGAM-SPP by eliminating the need for parameter tuning. Our modification, using the PLR (Hyvärinen and Smith, 2013) in place of the $k$NN-based mutual information estimators from earlier LiNGAM-SPP studies (Suzuki and Yang, 2022, 2024), not only eliminated the need for fine-tuning but also demonstrated superior performance and computational efficiency on simulated data. Furthermore, the modified LiNGAM-SPP we propose, PLR LiNGAM-SPP, proves to be more stable than DirectLiNGAM, significantly outperforming it specially in simulations with more features. Our method performs at least as well as DirectLiNGAM on real-world data and, in some cases, outperforms it by capturing more required edges and fewer forbidden edges.
Additionally, we introduced a functionality to LiNGAM-SPP for incorporating prior knowledge. Notably, our approach allows for specifying relative orderings rather than exact edges, offering increased flexibility. The integration of prior knowledge improved the method’s performance and reduced the number of edges traversed.
Furthermore, we extended the LiNGAM-SPP method by leveraging the entire path distribution. Specifically, the standardized moments of the path distribution were used as features to create models that estimate certain properties of the causal graph. We created a confounder detector, a sparsity estimator, and predictors for DirectLiNGAM and PLR LiNGAM-SPP performance. The confounder detector and sparsity estimator showed robust performance, achieving high AUCs of 0.783 and 0.774 respectively. Meanwhile, although the DirectLiNGAM and LiNGAM-SPP performance predictors achieved AUCs of 0.822 and 0.821 respectively, they exhibited a higher rate of false positives due to the inherent challenge in getting the exact causal ordering.
We also acknowledge the limitations of our study, such as the intractability of path enumeration for large $p$, which led us to limit our demonstration to a small number of features. This limitation is somewhat addressed using the sampling feature of ZDDs, but constructing the ZDD itself becomes more difficult (requiring more memory) as $p$ increases.
In conclusion, our study proposed an enhancement to the current LiNGAM-SPP and demonstrated other ways to leverage the path search formulation of the causal ordering problem. While further research is needed to address limitations and provide theoretical grounding, our results offer a promising direction for inferring properties of the causal graph from observational data.

### v2

@@ 3.2. PLR as an Independence Measure
Furthermore, Table 2 presents the average run times for the same cases. As expected, DirectLiNGAM, being a greedy algorithm, exhibits the fastest runtimes. Meanwhile, the runtimes of copent and $k$NN LiNGAM-SPP, both of which rely on the $k$NN mutual information estimator, tend to increase with sample size $N$. Despite this trend, copent LiNGAM-SPP has the smallest runtimes out of all LiNGAM-SPP methods. On the other hand, $k$NN LiNGAM-SPP runs notably slower. For instance, at $p=15$, the runtime can extend to hours. While copent LiNGAM-SPP is faster, its performance remains subpar. Overall, PLR LiNGAM-SPP presents the most favorable compromise between runtime and performance, with the added advantage of not requiring parameter tuning.
3.3. Testing on Real-World Data
To evaluate PLR LiNGAM-SPP on real-world data, we sourced datasets from the Carnegie Mellon University Causal Learning and Reasoning Group list of benchmarks accessible on GitHub (https://github.com/cmu-phil/example-causal-datasets). Our criteria for selection included datasets with continuous features and known ground truth. Additionally, we restricted our choices to datasets with $p\leq 20$ due to the computational demands of the LiNGAM-SPP method for larger values of $p$.
To derive the causal graph from LiNGAM-SPP, we adapted code from the lingam package (Ikeuchi et al., 2023). Specifically, after getting the causal order, we used the same sparse regression as DirectLiNGAM to identify which edges are zero (Zou, 2006; Shimizu et al., 2011). Table 3 provides the ground truth for each dataset, indicating required and forbidden edges, along with the edges captured by PLR DirectLiNGAM and PLR LiNGAM-SPP. The ground truth originally included tiers, which we converted into forbidden edges given that later tiers cannot cause earlier ones. For more details about the datasets and the provided ground truth, please refer to the GitHub repository.
Table 3. Performance of PLR DirectLiNGAM and PLR LiNGAM-SPP on Real-World Datasets
Known Edges DirectLiNGAM LiNGAM-SPP
@@ 3.3. Testing on Real-World Data
The results in Table 3 reveal that for the wine-quality-red and wine-quality-white datasets, PLR LiNGAM-SPP outperforms PLR DirectLiNGAM, capturing fewer forbidden edges and more required edges. For the other datasets, both methods yield identical results. This observation aligns with the findings in Table 1, indicating that PLR LiNGAM-SPP is more robust for longer sequences. This also suggests that PLR LiNGAM-SPP performs at least as well as PLR DirectLiNGAM in most cases.
4. Incorporating Known Relative Ordering
One notable feature of DirectLiNGAM is its capability to incorporate prior knowledge to enhance overall results by allowing users to specify which directed edges are required, forbidden, or unknown (Shimizu et al., 2011). In this study, we introduce an enhancement to LiNGAM-SPP to allow for integrating such prior knowledge. Unlike DirectLiNGAM, where the exact edges need to be specified, our implementation only requires specifying the relative order. In practical terms, this means we can assert that one feature causes another while still allowing for the possibility of having intermediary features, or mediators, between them.
To implement this, we modify the shortest path algorithm to selectively skip nodes that violate the specified order. The path formulation of LiNGAM-SPP naturally lends itself to this task. Each node in the graph representation of causal orderings contains information about relative orderings, indicating which features have been selected and which are yet to be explored. Specifically, if a feature is absent from the subscripts, it is positioned earlier in the causal order. For instance, referencing the graph in Figure 1, if we assert that $x_{1}$ precedes $x_{2}$, then nodes with subscript 11{1}1 but not 22{2}2 – such as $\{r_{3}^{(2)},r_{1}^{(2)}\}$ and $\{r_{1}^{(2,3)}\}$ – should be skipped. This is because they imply that $x_{2}$ comes before $x_{1}$, violating the known ordering.
We demonstrate the impact of incorporating prior knowledge on performance in Figure 2 and Table 4. To quantify the extent of prior knowledge, we express it as a percentage of variables with specified orderings. For example, if $p=8$, 50% knowledge means that the relative ordering of 4 variables is given. In this demonstration, the input is the relative ordering of 2 or more variables, which we then convert to variable pairs. As illustration, $(1,2,3)$ implies $(1,2)$, $(2,3)$, $(1,3)$. It is, however, also possible to manually specify pairs. As expected, incorporating more prior knowledge improves performance and generally reduces the traversed number of edges.
@@ 4. Incorporating Known Relative Ordering
Relative orderings provide a more flexible way of incorporating prior knowledge, which is valuable because we do not always know the exact edges. It is more common for us to know the relative ordering rather than the exact edges. For instance, in most applications, it is known that nothing can affect the age, sex, and other background variables of a person so these must be positioned earlier in the ordering relative to other variables. Another example is incorporating temporal information, where features with time-related attributes may exist and we only know the temporal order but not necessarily the ordering among the features within the same time slice.
5. Predicting Causal Graph Properties
In the preceding sections, the advantage of LiNGAM-SPP, which reframes the causal ordering problem as a shortest path problem, combined with original enhancements were discussed. However, we find that the utility of this formulation of LiNGAM extends beyond finding the shortest path. We discover that generating the distribution of the total path lengths from all possible orderings through path enumeration reveals crucial insights into the causal graph properties. Particularly, we observe that the path distribution may be used to infer graph properties such as the presence of unmeasured confounders, the sparsity of the graph, and to some extent, the reliability of PLR DirectLiNGAM and PLR LiNGAM-SPP. The presence of unmeasured confounders, to emphasize, is an essential assumption to verify especially for LiNGAM.
Our inspiration stems from the observations in Figure 3. These histograms illustrate the combined log-transformed PLR path distributions for $p=4,5,6$ using the same data generating process and simulation configurations as the previous sections. Notably, clear patterns emerge in the path distributions for the different cases. More challenging scenarios such as the presence of confounders or increased graph sparsity show increased frequency of shorter paths. This may indicate that correct ordering becomes more challenging and ambiguous in these cases.
@@ 5. Predicting Causal Graph Properties
To take advantage of the insight that the path distribution informs us about graph properties, we extract standardized moments to capture the shape of the distribution. Specifically, we compute the standardized moments of the log-transformed path lengths. These standardized moments will then serve as features in models trained to predict the various graph properties. The standardized moment is defined as
(14) $\tilde{\mu}_{k}=\frac{\text{E}[(X-\mu)^{k}]}{(\text{Var}[X-\mu])^{k/2}}.$
Here, $\tilde{\mu}_{k}$ denotes the $k$-th standardized moment and $X$ represents the log-transformed path lengths. To test the generalizability of the models, $p=4,5,6$ were used for training while $p=7$, unseen in the training set, is used for testing.
In the subsequent sections, we form a suite of classical machine learning algorithms. The $k$NN (Pedregosa et al., 2011), Random Forest (Pedregosa et al., 2011), XGBoost (Chen and Guestrin, 2016), CatBoost (Dorogush et al., 2018), and AdaBoost (Pedregosa et al., 2011)) are used to train the following models.
  1. (1)
    Confounder Detector – a classifier to detect the presence of unmeasured confounders.
  2. (2)
    Sparseness Estimator – a classifier and a regression model to estimate graph sparseness.
  3. (3)
    Performance Predictors – classifier and regression models to predict the reliability of causal discovery algorithms.
We use the receiver operating characteristic (ROC) curve as basis for the metric of discrimination, the area under the ROC curve (AUC).
5.1. Path Enumeration via ZDDs
Before delving into the details of the models, this subsection highlights our decision to use ZDDs for path enumeration. ZDDs are compact data structures that efficiently represent a family of sparse subsets from a finite ordered universe $U$ (Toda et al., 2017; Minato, 1993). In the context of graph theory, the diagram may also represent a family of subgraphs $G^{\prime}\subseteq G$ for a graph $G=(V,E)$. The frontier-based search algorithm provides a way to construct the diagram representing subgraphs, particularly paths, in a graph by setting the universe as the set of edges $U=E$ (Kawahara et al., 2017).
While the TdZdd library is commonly used for ZDD construction and set operations, its C++ implementation posed compatibility issues with our Python pipeline. Hence, we adopted the graphillion library instead, which offers a Python interface for ZDD operations. Although the standardized moments were calculated post hoc in our implementation, direct computation using ZDD operations, as demonstrated by (Lim et al., 2024), is also possible.
5.2. Confounder Detector
Detecting unmeasured confounders is crucial especially for LiNGAM, which assumes the absence of unmeasured confounders. In this section, we develop a classifier that detects the presence of unmeasured confounders. The features used were the 3rd to the 30th standard moments of the log-transformed path distribution and the target is binary – 1 if there are confounders and 0 if none. We then train different classifier models as shown in Table 5. Note that for demonstration purposes, these models were not fine-tuned, hence there is room for improving these results. Regardless, the models demonstrate good performance, with CatBoost achieving the highest AUC of $0.78$ that indicates good discrimination between the classes.
Table 5. Confounder Detectors Trained with Various Classifiers
@@ 5.2. Confounder Detector
Figure 4 shows the ROC curve and the confusion matrix from CatBoost evaluated on the test set. The optimal threshold is determined using Youden’s J statistic, which selects the threshold that maximizes the difference between the true positive and false positive rates. This threshold can be changed depending on the precision and sensitivity requirements of the use case. For the optimal threshold, we nevertheless see that the CatBoost model exhibits $0.77$ recall and $0.70$ precision.
@@
The figure illustrates the Receiver Operating Characteristic (ROC) curve and confusion matrix for the confounder detector model. Employing CatBoost and assessed on the test set, the optimal threshold, identified through Youden’s J statistic, maximizes the difference between true positive and false positive rates. The associated confusion matrix is derived from this threshold, showcasing the CatBoost model’s remarkable statistics ($0.77$ recall and $0.70$ precision).
We have thus shown that it is possible to detect unmeasured confounders with reasonable performance using information from the shape of the path distribution captured through the standardized moments.
5.3. Sparsity Estimator
Another task we explore is predicting graph sparsity using standard moments. Using the same features as the confounder detector, we initially create a classifier predicting whether sparsity is greater than $0.5$. Table 6 presents the results from different models. Similar to the confounder detector, CatBoost achieved the highest AUC of $0.77$. While the precision and recall of CatBoost are remarkable, AdaBoost offers slightly better precision and XGBoost provides slightly better recall. Choosing between them depends on the specific use case. Figure 5 displays the ROC curve and confusion matrix for CatBoost. Once again, it demonstrates good AUC, precision, and recall, indicating effective discrimination.
Table 6. Sparsity Estimators Trained with Various Classifiers
@@ 5.3. Sparsity Estimator
The figure illustrates the Receiver Operating Characteristic (ROC) curve and confusion matrix for the sparsity estimator model. Employing CatBoost and assessed on the test set, the optimal threshold, identified through Youden’s J statistic, maximizes the difference between true positive and false positive rates. The associated confusion matrix is derived from this threshold, showcasing the CatBoost model’s remarkable statistics ($0.72$ recall and $0.70$ precision).
In addition to the classifier, we explored creating a regression model using the same features with the target being the actual sparsity ranging from 0 to 1. Table 7 compares results from different models. Among them, AdaBoost delivered the best coefficient of determination $R^{2}$ of $0.16$, indicating that the shape of the distribution may explain some of the variance in sparsity. AdaBoost also yields a root-mean-square error (RMSE) of $0.27$, reasonably estimating sparsity.
Table 7. Regression Results for Sparsity Estimation
@@
5.4. Performance Predictors
So far, we have created models for detecting confounders and graph sparsity. These two properties are crucial for evaluating the performance of DirectLiNGAM and LiNGAM-SPP, as cases involving confounders or sparse graphs tend to pose greater challenges and lead to the poor performance of these methods. Further, we want to test if it is possible to directly predict the performance of these methods. In line with this, we develop classifiers to predict whether PLR DirectLiNGAM and PLR LiNGAM-SPP would produce the correct causal ordering. In other words, the target is 1 if the predicted causal order is entirely correct. Table 8 showcases the results.
Table 8. Performance Predictors Trained with Various Classifiers
@@ 5.4. Performance Predictors
For the LiNGAM-SPP performance predictor, CatBoost delivered the best AUC of 0.82. Given the relatively small number of instances with exact correct ordering, it is more likely to get some of the ordering wrong than to get everything right. Recall is therefore reasonable at $0.70$ but precision is not very high at $0.33$, indicating a significant number of false positives.
Meanwhile, for the DirectLiNGAM performance predictor, both the Random Forest and CatBoost provided similar AUC metrics of $0.82$. The Random Forest has the best recall at $0.80$ while another model, XGBoost, exhibits the highest precision at $0.31$. Similar to the LiNGAM-SPP case, the sensitivity of the models is good but precision is not as high, suggesting a notable presence of false positives.
We also explored fitting regression models to the ordering score of LiNGAM-SPP and DirectLiNGAM, with the results shown in Table 9. For the LiNGAM-SPP case, AdaBoost emerged as the best model, achieving a $0.21$ $R^{2}$ that indicates its ability to explain some variance in the ordering score. AdaBoost also boasted the lowest RMSE of $0.18$, reasonably pinpointing the ordering score within a reasonable range. For DirectLiNGAM, AdaBoost also outperformed other models, achieving an $R^{2}$ of $0.21$ and an RMSE of $0.18$.
@@
5.5. Remarks on Path Sampling
One drawback of path enumeration is the factorial increase of the number of edges with the number of features. To address the impracticality of using the entire path distribution especially in cases with large $p$, we opt for a more practical solution—sampling from the path distribution to compute the standardized moments. As mentioned earlier, ZDDs provide this functionality by sampling from the subset of graphs it represents. Here, we demonstrate the trade-off between path sample size and model performance, revealing that sufficiently good performance for the models discussed earlier may still be achieved without resorting to the entire path distribution.
Table 10 presents the performance of the models under various path sampling sizes. In this demonstration, we limit the training and testing to $p=7$ since $7!=5040$, which is sufficiently large for our evaluation. The table illustrates that the AUC of the model generally increases with the path sample size, approaching the performance of the full path distribution. For instance, even at a path sample size of 2500, the confounder detector and sparsity estimator already achieved AUCs that are sufficiently close to those obtained using the full path distribution. While the performance predictors for the same path sample size still have a slightly larger gap, the performance is acceptable. For larger values of $p$, even a fraction of the possible paths may already yield satisfactory performance.
@@ 6. Conclusion
We address a critical limitation of LiNGAM-SPP by eliminating the need for parameter tuning. Our modification, which uses the PLR (Hyvärinen and Smith, 2013) in place of the $k$NN-based mutual information estimators from earlier LiNGAM-SPP studies (Suzuki and Yang, 2022, 2024), not only eliminated the need for fine-tuning but also demonstrated superior performance and computational efficiency on simulated data. Furthermore, the modified LiNGAM-SPP we propose, PLR LiNGAM-SPP, proves to be more stable than DirectLiNGAM, significantly outperforming it in simulations with more features. Our method performs at least as well as DirectLiNGAM on real-world data and in some cases, outperforms it by capturing more required edges and fewer forbidden edges.
We also introduce a functionality to LiNGAM-SPP for incorporating prior knowledge. Our approach allows for specifying relative orderings rather than exact edges, offering increased flexibility. The integration of prior knowledge improved performance and reduced the number of edges traversed.
Furthermore, we extend the LiNGAM-SPP method by leveraging the entire path distribution. The standardized moments of the path distribution were used as features to create models that estimate certain properties of the true causal graph. We created a confounder detector, a sparsity estimator, and predictors for DirectLiNGAM and PLR LiNGAM-SPP performance. The confounder detector and sparsity estimator showed robust performance, achieving high AUCs of $0.78$ and $0.77$, respectively. Although the DirectLiNGAM and LiNGAM-SPP performance predictors achieved AUCs of $0.822$ and $0.821$, respectively, they exhibited a higher rate of false positives due to the inherent challenge in getting the exact causal ordering.
We acknowledge the limitations of our study, such as the intractability of path enumeration for large $p$, which led us to limit our demonstration to a small number of features. This limitation is somewhat addressed using the sampling feature of ZDDs, but constructing the ZDD itself becomes more difficult and requires more memory as $p$ increases.
To close, our study proposes enhancements to the current LiNGAM-SPP and demonstrates other ways to leverage the shortest path problem formulation of causal discovery frameworks. While further research is needed to address the limitations and provide theoretical grounding, our results offer a promising direction for inferring notable properties of the true causal graph from observational data.